1.         What is the Kb for B-, if the Ka for the acid (HB) is 2.5 x 10-8?

2.         What is the hydronium ion concentration of a 3.50 x 10-2 M solution of the monoprotic acid, HBrC6H4COO, for which Ka = 1.00 x 10-4?

3.         What is the Ka value for HB, if a solution that is 0.25 M B- and 0.55 M HB produces a pH of 4.95?

4.         Calculate the pH of a solution containing .225 moles of lactic acid (HC3H5O3) in water to produce one liter of solution. (Ka = 1.4 x 10-4)
Answer #1

(1) The relation between Kb , Ka & Kw is Kw = Ka x Kb

                                                          Kb = Kw / Ka

                                                              = (1.0x10-14) / (2.5x10-8)

                                                              = 4.0x10-7

Therefore Kb for B- 4.0x10-7

(2) Let a be the dissociation of the weak acid
                            HA <---> H + + A-

initial conc.            c               0         0

change                -ca            +ca      +ca

Equb. conc.         c(1-a)          ca       ca

Dissociation constant , Ka = ca x ca / ( c(1-a)

                                         = c a2 / (1-a)

In the case of weak acids a is very small so 1-a is taken as 1

So Ka = ca2

==> a = √ ( Ka / c )

Given Ka = 1.00x10-4

          c = concentration = 3.50x10-2 M

Plug the values we get a = 0.0534

So [H+] = ca = 3.50x10-2 M x 0.0534 = 1.87x10-3 M

Therefore the [H+] = 1.87x10-3 M

(3) According to Henderson's equation, pH = pKa + log ([B-]/[HB])

                                                       pKa = pH - log ([B-]/[HB])

                                                             = 4.95 - log(0.25/0.55)

                                                             = 5.29

(4) Let a be the dissociation of the weak acid
                            HA <---> H + + A-

initial conc.            c               0         0

change                -ca            +ca      +ca

Equb. conc.         c(1-a)          ca       ca

Dissociation constant , Ka = ca x ca / ( c(1-a)

                                         = c a2 / (1-a)

In the case of weak acids a is very small so 1-a is taken as 1

So Ka = ca2

==> a = √ ( Ka / c )

Given Ka = 1.4x10-4

          c = concentration = number of moles / volume in L

             = 0.225 mol / 1.0 L

              = 0.225 M

Plug the values we get a = 0.0249

[H+] = ca = 0.225 M x 0.0249 = 5.612x10-3 M

pH = - log[H+] = 2.25